# Robust Bayesian Inference for Berkson and Classical Measurement Error Models

### v2

@@ B.2 Proofs of theoretical results for the Classical error model
$\quad\quad\quad\frac{1}{\sqrt{n}(c+1)}+\frac{1}{\sqrt{c+2}}+\frac{c}{c+1}\operatorname{\operatorname{MMD}}_{k_{X}}(\operatorname{\mathbb{P}}_{X}^{0},\operatorname{\mathbb{F}})+\frac{1}{c+1}\operatorname{\operatorname{MMD}}_{k_{X}}(\operatorname{\mathbb{F}}_{N}^{0},\delta_{0})$
@@
We use the stick-breaking process representation (Sethuraman,, 1994) of a DP, meaning that for each $i\in\{1,\dots,n\}$, $\operatorname{\mathbb{P}}^{i}=\sum_{j=1}^{\infty}\xi_{j}^{i}\delta_{z_{j}^{i}}$ where $\xi_{1:\infty}\sim\operatorname{\operatorname{GEM}}(c+1)$ and $z_{1:\infty}\operatorname{\overset{iid}{\sim}}\operatorname{\mathbb{F}}^{\prime}_{w_{i}}$. For clarity we use the separate expectations $\operatorname{\mathbb{E}}_{\xi^{i}_{1:\infty}\sim\operatorname{\operatorname{GEM}}(c+1)}$ and $\mathbb{E}_{z^{i}_{1:\infty}\operatorname{\overset{iid}{\sim}}\operatorname{\mathbb{F}}^{\prime}_{w_{i}}}$ induced by $\mu_{i}$ for each $i\in\{1,\dots,n\}$. Further let $\operatorname{\mathbb{F}}^{\prime}:=\frac{1}{n}\sum_{i=1}^{n}\operatorname{\mathbb{F}}^{\prime}_{w_{i}}$. Using the triangle inequality we have
$\operatorname{\operatorname{MMD}}_{k_{X}}(\mathbb{P}_{X}^{0},\mathbb{P}_{X})\leq\operatorname{\operatorname{MMD}}_{k_{X}}(\operatorname{\mathbb{P}}_{X},\operatorname{\mathbb{F}}^{\prime})+\operatorname{\operatorname{MMD}}_{k_{X}}(\mathbb{P}_{X}^{0},\operatorname{\mathbb{F}}^{\prime})$ (40)
$\mathbb{E}_{\operatorname{\mathbb{P}}\sim\mu}[\operatorname{\operatorname{MMD}}^{2}_{k_{X}}(\operatorname{\mathbb{P}}_{X},\operatorname{\mathbb{F}}^{\prime})]$ $=\mathbb{E}_{\operatorname{\mathbb{P}}\sim\mu}[\|\mu_{\operatorname{\mathbb{P}}_{X}}-\mu_{\operatorname{\mathbb{F}}^{\prime}}\|_{\mathcal{H}_{k_{X}}}^{2}]$
$=\mathbb{E}_{\operatorname{\mathbb{P}}\sim\mu}\left[\left\|\frac{1}{n}\sum_{i=1}^{n}(\mu_{\operatorname{\mathbb{P}}^{i}}-\mu_{\operatorname{\mathbb{F}}^{\prime}_{w_{i}}})\right\|^{2}_{\mathcal{H}_{k_{X}}}\right]$
$=\operatorname{\mathbb{E}}_{\xi^{1:n}_{1:\infty}\sim\operatorname{\operatorname{GEM}}(c+1)}\left[\mathbb{E}_{z^{1:n}_{1:\infty}\operatorname{\overset{iid}{\sim}}\operatorname{\mathbb{F}}^{\prime}_{w_{i}}}\left[\left\|\frac{1}{n}\sum_{i=1}^{n}\left(\sum_{j=1}^{\infty}\xi_{j}^{i}k_{X}(z_{j}^{i},\cdot)-\mu_{\operatorname{\mathbb{F}}^{\prime}_{w_{i}}}\right)\right\|^{2}_{\mathcal{H}_{k_{X}}}\right]\right]$
$\leq\frac{1}{n}\sum_{i=1}^{n}\operatorname{\mathbb{E}}_{\xi^{i}_{1:\infty}\sim\operatorname{\operatorname{GEM}}(c+1)}\left[\mathbb{E}_{z^{i}_{1:\infty}\operatorname{\overset{iid}{\sim}}\operatorname{\mathbb{F}}^{\prime}_{w_{i}}}\left[\left\|\sum_{j=1}^{\infty}\xi_{j}^{i}k_{X}(z_{j}^{i},\cdot)-\mu_{\operatorname{\mathbb{F}}^{\prime}_{w_{i}}}\right\|^{2}_{\mathcal{H}_{k_{X}}}\right]\right]$
$=\frac{1}{n}\sum_{i=1}^{n}\operatorname{\mathbb{E}}_{\xi^{i}_{1:\infty}\sim\operatorname{\operatorname{GEM}}(c+1)}\left[\mathbb{E}_{z^{i}_{1:\infty}\operatorname{\overset{iid}{\sim}}\operatorname{\mathbb{F}}^{\prime}_{w_{i}}}\left[\left\|\sum_{j=1}^{\infty}\xi_{j}^{i}\left(k_{X}(z_{j}^{i},\cdot)-\mu_{\operatorname{\mathbb{F}}^{\prime}_{w_{i}}}\right)\right\|^{2}_{\mathcal{H}_{k_{X}}}\right]\right]$
$=\frac{1}{n}\sum_{i=1}^{n}\operatorname{\mathbb{E}}_{\xi^{i}_{1:\infty}\sim\operatorname{\operatorname{GEM}}(c+1)}\left[\mathbb{E}_{z^{i}_{1:\infty}\operatorname{\overset{iid}{\sim}}\operatorname{\mathbb{F}}^{\prime}_{w_{i}}}\left[\sum_{j=1}^{\infty}(\xi_{j}^{i})^{2}\left\|k_{X}(z_{j}^{i},\cdot)-\mu_{\operatorname{\mathbb{F}}^{\prime}_{w_{i}}}\right\|^{2}_{\mathcal{H}_{k_{X}}}\right.\right.$
$\quad\quad\left.\left.+2\sum_{j\neq t}\xi_{j}^{i}\xi_{t}^{i}\left<k_{X}(z_{j},\cdot)-\mu_{\operatorname{\mathbb{F}}^{\prime}_{w_{i}}},k_{X}(z_{t},\cdot)-\mu_{\operatorname{\mathbb{F}}^{\prime}_{w_{i}}}\right>_{\operatorname{\mathcal{H}_{k_{X}}}}\right]\right].$ (41)
For any $j\neq t$ and each $i\in\{1,\dots,n\}$ we have:
$=\operatorname{\mathbb{E}}_{z_{j},z_{t}\operatorname{\overset{iid}{\sim}}\operatorname{\mathbb{F}}^{\prime}_{w_{i}}}\left[\left<k_{X}(z_{j},\cdot)-\operatorname{\mathbb{E}}_{z_{j}\sim\operatorname{\mathbb{F}}^{\prime}_{w_{i}}}[k_{X}(z_{j},\cdot)],k_{X}(z_{t},\cdot)-\operatorname{\mathbb{E}}_{z_{t}\sim\operatorname{\mathbb{F}}^{\prime}}[(k_{X}(z_{t},\cdot)]\right>_{\operatorname{\mathcal{H}_{k_{X}}}}\right]$
$\qquad-\operatorname{\mathbb{E}}_{z_{t}\sim\operatorname{\mathbb{F}}_{w_{i}}}[k_{X}(z_{t},\cdot)]\operatorname{\mathbb{E}}_{z_{j}\sim\operatorname{\mathbb{F}}_{w_{i}}}[k_{X}(z_{j},\cdot)]+\operatorname{\mathbb{E}}_{z_{t}\sim\operatorname{\mathbb{F}}^{\prime}_{w_{i}}}[k_{X}(z_{t},\cdot)]\operatorname{\mathbb{E}}_{z_{j}\sim\operatorname{\mathbb{F}}^{\prime}_{w_{i}}}[k_{X}(z_{j},\cdot)]$
$=\operatorname{\mathbb{E}}_{z_{j},z_{t}\operatorname{\overset{iid}{\sim}}\operatorname{\mathbb{F}}^{\prime}_{w_{i}}}[k_{X}(z_{j},z_{t})]-\operatorname{\mathbb{E}}_{z_{j},z_{t}\operatorname{\overset{iid}{\sim}}\operatorname{\mathbb{F}}^{\prime}_{w_{i}}}[k_{X}(z_{j},z_{t})]-\operatorname{\mathbb{E}}_{z_{j},z_{t}\operatorname{\overset{iid}{\sim}}\operatorname{\mathbb{F}}^{\prime}_{w_{i}}}[k_{X}(z_{j},z_{t})]+\operatorname{\mathbb{E}}_{z_{j},z_{t}\operatorname{\overset{iid}{\sim}}\operatorname{\mathbb{F}}^{\prime}_{w_{i}}}[k_{X}(z_{j},z_{t})]$
$=0.$ (42)
Moreover, for any $j=1,2,\dots$ and each $i\in\{1,\dots,n\}$ we have
$\operatorname{\mathbb{E}}_{z_{k}\sim\operatorname{\mathbb{F}}^{\prime}_{w_{i}}}\left[\|k_{X}(z_{k},\cdot)-\mu_{\operatorname{\mathbb{F}}^{\prime}_{w_{i}}}\|_{\operatorname{\mathcal{H}_{k_{X}}}}^{2}\right]$ $=\operatorname{\mathbb{E}}_{z_{k}\sim\operatorname{\mathbb{F}}^{\prime}_{w_{i}}}\left[\|k_{X}(z_{k},\cdot)-\operatorname{\mathbb{E}}_{z_{k}\sim\operatorname{\mathbb{F}}^{\prime}_{w_{i}}}[k_{X}(z_{k},\cdot)]\|_{\operatorname{\mathcal{H}_{k_{X}}}}^{2}\right]$
$=\operatorname{\mathbb{E}}_{z_{k}\sim\operatorname{\mathbb{F}}^{\prime}_{w_{i}}}\left[\|k_{X}(z_{k},\cdot)\|_{\operatorname{\mathcal{H}_{k_{X}}}}^{2}-2\left<k_{X}(z_{k},\cdot),\operatorname{\mathbb{E}}_{z_{k}\sim\operatorname{\mathbb{F}}_{w_{i}}}[k_{X}(z_{k},\cdot)]\right>\right.$
$\quad\quad\left.+\|\operatorname{\mathbb{E}}_{z_{k}\sim\operatorname{\mathbb{F}}^{\prime}_{w_{i}}}[k_{X}(z_{k},\cdot)]\|_{\operatorname{\mathcal{H}_{k_{X}}}}^{2}\right]$
$=\operatorname{\mathbb{E}}_{z_{k}\sim\operatorname{\mathbb{F}}^{\prime}_{w_{i}}}\left[\|k_{X}(z_{k},\cdot)\|_{\operatorname{\mathcal{H}_{k_{X}}}}^{2}\right]-2\left<\operatorname{\mathbb{E}}_{z_{k}\sim\operatorname{\mathbb{F}}^{\prime}_{w_{i}}}[k_{X}(z_{k},\cdot)],\right.$
$\quad\quad\left.\operatorname{\mathbb{E}}_{z_{k}\sim\operatorname{\mathbb{F}}^{\prime}_{w_{i}}}[k_{X}(z_{k},\cdot)]\right>+\|\operatorname{\mathbb{E}}_{z_{k}\sim\operatorname{\mathbb{F}}^{\prime}_{w_{i}}}[k_{X}(z_{k},\cdot)]\|_{\operatorname{\mathcal{H}_{k_{X}}}}^{2}$
$=\operatorname{\mathbb{E}}_{z_{k}\sim\operatorname{\mathbb{F}}^{\prime}_{w_{i}}}\left[\|k(z_{k},\cdot)\|_{\operatorname{\mathcal{H}_{k_{X}}}}^{2}\right]-2\|\operatorname{\mathbb{E}}_{z_{k}\sim\operatorname{\mathbb{F}}^{\prime}_{w_{i}}}[k_{X}(z_{k},\cdot)]\|_{\operatorname{\mathcal{H}_{k_{X}}}}^{2}$
$\quad\quad+\|\operatorname{\mathbb{E}}_{z_{k}\sim\operatorname{\mathbb{F}}^{\prime}}[k_{X}(z_{k},\cdot)]\|_{\operatorname{\mathcal{H}_{k_{X}}}}^{2}$
$=\operatorname{\mathbb{E}}_{z_{k}\sim\operatorname{\mathbb{F}}^{\prime}_{w_{i}}}[\|k_{X}(z_{k},\cdot)\|_{\operatorname{\mathcal{H}_{k_{X}}}}^{2}]-\|\operatorname{\mathbb{E}}_{z_{k}\sim\operatorname{\mathbb{F}}^{\prime}_{w_{i}}}[k_{X}(z_{k},\cdot)]\|_{\operatorname{\mathcal{H}_{k_{X}}}}^{2}$
$\leq\operatorname{\mathbb{E}}_{z_{k}\sim\operatorname{\mathbb{F}}^{\prime}_{w_{i}}}[\|k_{X}(z_{k},\cdot)\|_{\operatorname{\mathcal{H}_{k_{X}}}}^{2}]$
$=\operatorname{\mathbb{E}}_{z_{k}\sim\operatorname{\mathbb{F}}^{\prime}_{w_{i}}}[|k_{X}(z_{k},z_{k})|].$ (43)
Substituting (42) and (43) in (41):
$\mathbb{E}_{\operatorname{\mathbb{P}}\sim\mu}[\operatorname{\operatorname{MMD}}^{2}_{k_{X}}(\operatorname{\mathbb{P}}_{X},\operatorname{\mathbb{F}}^{\prime})]$ $\leq\frac{1}{n}\sum_{i=1}^{n}\sum_{j=1}^{\infty}\operatorname{\mathbb{E}}_{\xi_{1:\infty}^{i}\sim\operatorname{\operatorname{GEM}}(c+1)}[(\xi_{j}^{i})^{2}]\operatorname{\mathbb{E}}_{z_{k}\sim\operatorname{\mathbb{F}}^{\prime}_{w_{i}}}[|k_{X}(z_{k},z_{k})|]$
$\leq\frac{1}{n}\sum_{i=1}^{n}\sum_{j=1}^{\infty}\operatorname{\mathbb{E}}_{\xi_{1:\infty}^{i}\sim\operatorname{\operatorname{GEM}}(c+1)}[(\xi_{j}^{i})^{2}]$
$=\sum_{j=1}^{\infty}\frac{2(c+1)^{j-1}}{(c+3)^{j}(c+2)}$
$=\frac{2}{(c+3)(c+2)}\sum_{j=1}^{\infty}\left(\frac{c+1}{c+3}\right)^{j-1}$
$=\frac{2}{(c+3)(c+2)}\frac{1}{1-\frac{c+1}{c+3}}$
$=\frac{1}{c+2}$
$\mathbb{E}_{\operatorname{\mathbb{P}}\sim\mu}[\operatorname{\operatorname{MMD}}_{k_{X}}(\operatorname{\mathbb{P}}_{X},\operatorname{\mathbb{F}}^{\prime})]\leq\sqrt{\mathbb{E}_{\operatorname{\mathbb{P}}\sim\mu}[\operatorname{\operatorname{MMD}}^{2}_{k_{X}}(\operatorname{\mathbb{P}}_{X},\operatorname{\mathbb{F}}^{\prime})]}\leq\frac{1}{\sqrt{c+2}}.$ (44)
@@
$\mu_{\operatorname{\mathbb{F}}^{\prime}}:=\mathbb{E}_{z\sim\operatorname{\mathbb{F}}^{\prime}}[k(z,\cdot)]=\frac{c}{c+1}\operatorname{\mathbb{E}}_{z\sim\operatorname{\mathbb{F}}}[k(z,\cdot)]+\frac{1}{c+1}\mathbb{E}_{z\sim\operatorname{\mathbb{P}}_{W}^{n}}[k(z,\cdot)]=\frac{c}{c+1}\mu_{\operatorname{\mathbb{F}}}+\frac{1}{c+1}\mu_{\operatorname{\mathbb{P}}_{W}^{n}}.$
@@
$\operatorname{\operatorname{MMD}}_{k_{X}}(\operatorname{\mathbb{P}}_{X}^{0},\operatorname{\mathbb{F}}^{\prime})$ $=\left\|\mu_{\operatorname{\mathbb{P}}_{X}^{0}}-\frac{c}{c+1}\mu_{\operatorname{\mathbb{F}}}-\frac{1}{c+1}\mu_{\operatorname{\mathbb{P}}_{W}^{n}}\right\|_{\operatorname{\mathcal{H}_{k_{X}}}}$
$\leq\frac{c}{c+1}\|\mu_{\operatorname{\mathbb{P}}_{X}^{0}}-\mu_{\operatorname{\mathbb{F}}}\|_{\operatorname{\mathcal{H}_{k_{X}}}}+\frac{1}{c+1}\|\mu_{\operatorname{\mathbb{P}}_{X}^{0}}-\mu_{\operatorname{\mathbb{P}}_{W}^{n}}\|_{\operatorname{\mathcal{H}_{k_{X}}}}$
$\leq\frac{c}{c+1}\|\mu_{\operatorname{\mathbb{P}}_{X}^{0}}-\mu_{\operatorname{\mathbb{F}}}\|_{\operatorname{\mathcal{H}_{k_{X}}}}+\frac{1}{c+1}\|\mu_{\operatorname{\mathbb{P}}_{X}^{n}}-\mu_{\operatorname{\mathbb{P}}_{W}^{n}}\|_{\operatorname{\mathcal{H}_{k_{X}}}}+\frac{1}{c+1}\|\mu_{\operatorname{\mathbb{P}}_{X}^{0}}-\mu_{\operatorname{\mathbb{P}}_{X}^{n}}\|_{\operatorname{\mathcal{H}_{k_{X}}}}$
$=\frac{c}{c+1}\operatorname{\operatorname{MMD}}_{k_{X}}(\operatorname{\mathbb{P}}_{X}^{0},\operatorname{\mathbb{F}})+\frac{1}{c+1}\operatorname{\operatorname{MMD}}_{k_{X}}(\operatorname{\mathbb{P}}_{X}^{n},\operatorname{\mathbb{P}}_{W}^{n})+\frac{1}{c+1}\operatorname{\operatorname{MMD}}_{k_{X}}(\operatorname{\mathbb{P}}_{X}^{0},\operatorname{\mathbb{P}}_{X}^{n})$ (45)
where we used the triangle inequality for both inequalities. Using Lemma B.1 with $S_{i}=x_{i}\sim Q_{i}:=\mathbb{P}^{0}_{X}$ we have
@@
$\operatorname{\mathbb{E}}_{x_{1:n}\sim\operatorname{\mathbb{P}}^{0}_{X}}[\operatorname{\operatorname{MMD}}_{k_{X}}(\operatorname{\mathbb{P}}_{X}^{n},\operatorname{\mathbb{P}}_{W}^{n})]$ $=\operatorname{\mathbb{E}}_{x_{1:n}\sim\operatorname{\mathbb{P}}^{0}}\left[\left\|\frac{1}{n}\sum_{i=1}^{n}\left[\phi_{X}(x_{i})-\phi_{X}(w_{i})\right]\right\|_{\mathcal{H}_{k_{X}}}\right]$
$=\operatorname{\mathbb{E}}_{\nu^{0}_{1:n}\sim\operatorname{\mathbb{F}}_{N}^{0}}\left[\left\|\frac{1}{n}\sum_{i=1}^{n}\left[\phi_{X}(w_{i}-\nu_{i}^{0})-\phi_{X}(w_{i})\right]\right\|_{\mathcal{H}_{k_{X}}}\right]$
$=\left\|\frac{1}{n}\sum_{i=1}^{n}\left[\operatorname{\mathbb{E}}_{\nu^{0}_{i}\sim\operatorname{\mathbb{F}}_{N}^{0}}[k^{i}_{X}(\nu_{i}^{0},\cdot)]-k^{i}_{X}(0,\cdot)\right]\right\|_{\mathcal{H}_{k_{X}}}$
$\leq\frac{1}{n}\sum_{i=1}^{n}\operatorname{\operatorname{MMD}}_{k^{i}_{X}}(\operatorname{\mathbb{F}}_{N}^{0},\delta_{0})$
@@
$\mathbb{E}_{x_{1:n}\sim\operatorname{\mathbb{P}}_{X}^{0}}[\operatorname{\operatorname{MMD}}_{k_{X}}(\operatorname{\mathbb{P}}_{X}^{0},\operatorname{\mathbb{F}}^{\prime})]\leq\frac{1}{c+1}\frac{1}{\sqrt{n}}+\frac{c}{c+1}\operatorname{\operatorname{MMD}}_{k_{X}}(\operatorname{\mathbb{P}}_{X}^{0},\operatorname{\mathbb{F}})+\frac{1}{c+1}\operatorname{\operatorname{MMD}}_{k_{X}}(\operatorname{\mathbb{F}}_{N}^{0},\delta_{0}).$ (46)
Finally, using (44) and (46) in (40) we obtain the required result:
@@
$\quad\quad\quad\frac{1}{\sqrt{n}(c+1)}+\frac{1}{\sqrt{c+2}}+\frac{c}{c+1}\operatorname{\operatorname{MMD}}_{k_{X}}(\operatorname{\mathbb{P}}_{X}^{0},\operatorname{\mathbb{F}})+\frac{1}{c+1}\operatorname{\operatorname{MMD}}_{k_{X}}(\operatorname{\mathbb{F}}_{N}^{0},\delta_{0})$
Lemma B.7.
(Classical) Let $\mathbb{P}^{0,n}:=\operatorname{\mathbb{P}}_{X}^{n}\operatorname{\mathbb{P}}_{Y}^{n}=\frac{1}{n}\sum_{i=1}^{n}\delta_{(x_{i},y_{i})}$ and $\mathbb{P}_{XY}=\operatorname{\mathbb{P}}_{X}\operatorname{\mathbb{P}}_{Y}^{n}=\frac{1}{n}\sum_{i=1}^{n}\sum_{j=1}^{\infty}\xi_{j}^{i}\delta_{(z_{j},y_{i})}$ where $z^{i}_{1:\infty}\overset{iid}{\sim}\operatorname{\mathbb{F}}^{\prime}_{w_{i}},\xi^{i}_{1:\infty}\sim\operatorname{\operatorname{GEM}}(c+1)$. Then under assumptions 1-3:
$\quad\quad\quad\leq\frac{c+2}{\sqrt{n}(c+1)}+\frac{1}{\sqrt{c+2}}+\frac{c}{c+1}\operatorname{\operatorname{MMD}}_{k^{2}_{X}}(\operatorname{\mathbb{P}}_{X}^{0},\operatorname{\mathbb{F}})+\frac{1}{c+1}\operatorname{\operatorname{MMD}}_{k^{2}_{X}}(\operatorname{\mathbb{F}}_{N}^{0},\delta_{0})$

### v3

@@ D.2 Proofs of theoretical results for the Classical error model
$\quad\quad\quad\frac{c}{\sqrt{n}(c+1)}+\frac{1}{\sqrt{n(c+2)}}+\frac{c}{c+1}\operatorname{\operatorname{MMD}}_{k_{X}}(\operatorname{\mathbb{P}}_{X}^{0},\operatorname{\mathbb{F}})+\frac{1}{c+1}\operatorname{\operatorname{MMD}}_{k_{X}}(\operatorname{\mathbb{F}}_{N}^{0},\delta_{0})$
@@
Using the triangle inequality, we have
$\operatorname{\operatorname{MMD}}_{k_{X}}(\mathbb{P}_{X}^{0},\mathbb{P}_{X})\leq\operatorname{\operatorname{MMD}}_{k_{X}}(\operatorname{\mathbb{P}}_{X},\operatorname{\mathbb{F}}^{\prime})+\operatorname{\operatorname{MMD}}_{k_{X}}(\mathbb{P}_{X}^{0},\operatorname{\mathbb{F}}^{\prime}).$ (37)
The first term can be bounded exactly in the same way as the first part of the proof of Lemma D.2 since it does not depend on the error structure:
$\mathbb{E}_{x_{1:n}\sim\operatorname{\mathbb{P}}_{X}^{0}}[\operatorname{\mathbb{E}}_{\operatorname{\mathbb{P}}\sim\mu}[\operatorname{\operatorname{MMD}}_{k}(\operatorname{\mathbb{P}}_{X},\operatorname{\mathbb{F}}^{\prime})]]\leq\frac{1}{\sqrt{n(c+2)}}.$ (38)
@@
where $\operatorname{\mathbb{F}}:=\frac{1}{n}\sum_{i=1}^{n}\operatorname{\mathbb{F}}_{w_{i}}$ and $\operatorname{\mathbb{P}}_{W}:=\frac{1}{n}\sum_{i=1}\delta_{w_{i}}$ is the empirical measure of the observed, noisy covariates. By linearity of expectations it follows that
$\operatorname{\mathbb{E}}_{\operatorname{\mathbb{F}}^{\prime}}[\cdot]=\frac{c}{c+1}\operatorname{\mathbb{E}}_{\operatorname{\mathbb{F}}}[\cdot]+\frac{1}{c+1}\operatorname{\mathbb{E}}_{\operatorname{\mathbb{P}}_{W}^{n}}[\cdot]$
and the kernel mean embedding of $\operatorname{\mathbb{F}}^{\prime}$ is
$\mu_{\operatorname{\mathbb{F}}^{\prime}}$ $:=\mathbb{E}_{z\sim\operatorname{\mathbb{F}}^{\prime}}[k(z,\cdot)]=\frac{c}{c+1}\operatorname{\mathbb{E}}_{z\sim\operatorname{\mathbb{F}}}[k(z,\cdot)]+\frac{1}{c+1}\mathbb{E}_{z\sim\operatorname{\mathbb{P}}_{W}^{n}}[k(z,\cdot)]$
$=\frac{c}{c+1}\mu_{\operatorname{\mathbb{F}}}+\frac{1}{c+1}\mu_{\operatorname{\mathbb{P}}_{W}^{n}}.$
@@
$\operatorname{\operatorname{MMD}}_{k_{X}}(\operatorname{\mathbb{P}}_{X}^{0},\operatorname{\mathbb{F}}^{\prime})$
$\quad=\left\|\mu_{\operatorname{\mathbb{P}}_{X}^{0}}-\frac{c}{c+1}\mu_{\operatorname{\mathbb{F}}}-\frac{1}{c+1}\mu_{\operatorname{\mathbb{P}}_{W}^{n}}\right\|_{\operatorname{\mathcal{H}_{k_{X}}}}$
$\quad\leq\frac{c}{c+1}\|\mu_{\operatorname{\mathbb{P}}_{X}^{0}}-\mu_{\operatorname{\mathbb{F}}}\|_{\operatorname{\mathcal{H}_{k_{X}}}}+\frac{1}{c+1}\|\mu_{\operatorname{\mathbb{P}}_{X}^{0}}-\mu_{\operatorname{\mathbb{P}}_{W}^{n}}\|_{\operatorname{\mathcal{H}_{k_{X}}}}$
$\quad\leq\frac{c}{c+1}\|\mu_{\operatorname{\mathbb{P}}_{X}^{0}}-\mu_{\operatorname{\mathbb{F}}}\|_{\operatorname{\mathcal{H}_{k_{X}}}}+\frac{1}{c+1}\|\mu_{\operatorname{\mathbb{P}}_{X}^{n}}-\mu_{\operatorname{\mathbb{P}}_{W}^{n}}\|_{\operatorname{\mathcal{H}_{k_{X}}}}+\frac{1}{c+1}\|\mu_{\operatorname{\mathbb{P}}_{X}^{0}}-\mu_{\operatorname{\mathbb{P}}_{X}^{n}}\|_{\operatorname{\mathcal{H}_{k_{X}}}}$
$\quad=\frac{c}{c+1}\operatorname{\operatorname{MMD}}_{k_{X}}(\operatorname{\mathbb{P}}_{X}^{0},\operatorname{\mathbb{F}})+\frac{1}{c+1}\operatorname{\operatorname{MMD}}_{k_{X}}(\operatorname{\mathbb{P}}_{X}^{n},\operatorname{\mathbb{P}}_{W}^{n})+\frac{1}{c+1}\operatorname{\operatorname{MMD}}_{k_{X}}(\operatorname{\mathbb{P}}_{X}^{0},\operatorname{\mathbb{P}}_{X}^{n})$ (39)
where we used the triangle inequality for both inequalities. Using Lemma D.1 with $S_{i}=x_{i}\sim Q_{i}:=\mathbb{P}^{0}_{X}$ we have
@@
$\operatorname{\mathbb{E}}_{x_{1:n}\sim\operatorname{\mathbb{P}}^{0}_{X}}[\operatorname{\operatorname{MMD}}_{k_{X}}(\operatorname{\mathbb{P}}_{X}^{n},\operatorname{\mathbb{P}}_{W}^{n})]$
$\quad=\operatorname{\mathbb{E}}_{x_{1:n}\sim\operatorname{\mathbb{P}}^{0}}\left[\left\|\frac{1}{n}\sum_{i=1}^{n}\left[\phi_{X}(x_{i})-\phi_{X}(w_{i})\right]\right\|_{\mathcal{H}_{k_{X}}}\right]$
$\quad=\operatorname{\mathbb{E}}_{\nu^{0}_{1:n}\sim\operatorname{\mathbb{F}}_{N}^{0}}\left[\left\|\frac{1}{n}\sum_{i=1}^{n}\left[\phi_{X}(w_{i}-\nu_{i}^{0})-\phi_{X}(w_{i})\right]\right\|_{\mathcal{H}_{k_{X}}}\right]$
$\quad=\left\|\frac{1}{n}\sum_{i=1}^{n}\left[\operatorname{\mathbb{E}}_{\nu^{0}_{i}\sim\operatorname{\mathbb{F}}_{N}^{0}}[k^{i}_{X}(\nu_{i}^{0},\cdot)]-k^{i}_{X}(0,\cdot)\right]\right\|_{\mathcal{H}_{k_{X}}}$
$\quad\leq\frac{1}{n}\sum_{i=1}^{n}\operatorname{\operatorname{MMD}}_{k^{i}_{X}}(\operatorname{\mathbb{F}}_{N}^{0},\delta_{0})$
$\quad=\operatorname{\operatorname{MMD}}_{k_{X}}(\operatorname{\mathbb{F}}_{N}^{0},\delta_{0}).$
@@
$\mathbb{E}_{x_{1:n}\sim\operatorname{\mathbb{P}}_{X}^{0}}[\operatorname{\operatorname{MMD}}_{k_{X}}(\operatorname{\mathbb{P}}_{X}^{0},\operatorname{\mathbb{F}}^{\prime})]$ $\leq\frac{1}{c+1}\frac{1}{\sqrt{n}}+\frac{c}{c+1}\operatorname{\operatorname{MMD}}_{k_{X}}(\operatorname{\mathbb{P}}_{X}^{0},\operatorname{\mathbb{F}})$
$+\frac{1}{c+1}\operatorname{\operatorname{MMD}}_{k_{X}}(\operatorname{\mathbb{F}}_{N}^{0},\delta_{0}).$ (40)
Finally, using (38) and (D.2.1) in (37) we obtain the required result:
@@
$\quad\quad\quad\frac{c}{\sqrt{n}(c+1)}+\frac{1}{\sqrt{n(c+2)}}+\frac{c}{c+1}\operatorname{\operatorname{MMD}}_{k_{X}}(\operatorname{\mathbb{P}}_{X}^{0},\operatorname{\mathbb{F}})+\frac{1}{c+1}\operatorname{\operatorname{MMD}}_{k_{X}}(\operatorname{\mathbb{F}}_{N}^{0},\delta_{0})$
Lemma D.7.
(Classical) Let
$\mathbb{P}^{0,n}:=\frac{1}{n}\sum_{i=1}^{n}\delta_{(x_{i},y_{i})}\quad\text{and}\quad\mathbb{P}_{XY}=\operatorname{\mathbb{P}}_{X}\times\operatorname{\mathbb{P}}_{Y}^{n}=\frac{1}{n}\sum_{i=1}^{n}\sum_{j=1}^{\infty}\xi_{j}^{i}\delta_{(z_{j},y_{i})}$
where $z^{i}_{1:\infty}\overset{iid}{\sim}\operatorname{\mathbb{F}}^{\prime}_{w_{i}},\xi^{i}_{1:\infty}\sim\operatorname{\operatorname{GEM}}(c+1)$. Then under Assumptions 4.0.1-4.0.3:
$\mathbb{E}_{(x,y)_{1:n}\sim\operatorname{\mathbb{P}}^{0}}[\operatorname{\mathbb{E}}_{\operatorname{\mathbb{P}}\sim\mu}[\operatorname{\operatorname{MMD}}_{k}(\mathbb{P}^{0},\mathbb{P}_{XY})]]$
$\quad\leq\frac{1}{\sqrt{n(c+2)}}+\frac{c}{n(c+1)}\sum_{i=1}^{n}\sqrt{{\text{Var}}_{\mathcal{H}_{k_{X}}}[\operatorname{\mathbb{P}}_{X}^{0}]+\operatorname{\operatorname{MMD}}^{2}_{k_{X}}(\operatorname{\mathbb{P}}_{X}^{0},\operatorname{\mathbb{F}}_{X\mid w_{i}})}$
$\quad\quad+\frac{\sqrt{2}}{c+1}\sqrt{\operatorname{\operatorname{MMD}}_{k_{X}}(\operatorname{\mathbb{F}}_{N}^{0},\delta_{0})}.$